OpenStudy (anonymous):

Which of the following is the radical expression of 4d^3/8

OpenStudy (anonymous):

8root 4d^3 I don't know if that simplifies further.

OpenStudy (anonymous):

\[(4d)^{3/8} = ((4d)^{3})^{1/8} = \sqrt[8]{(4d)^{3}}\]